# Whom do Explanations Serve? A Systematic Literature Survey of User Characteristics in Explainable Recommender Systems Evaluation

### v1

@@ 5.1. Who was recruited to evaluate recommender systems explanations?
The results about the education level of the participants show that the majority of the participants on which explanations are evaluated are highly educated. Participants with at least a bachelor’s degree were the dominant group in our sample. The average participant has a higher education level than the average population.
@@ 5.3. Which user characteristics impact the effect of explanations in recommender systems?
In this survey, we looked at four potential factors, the explanation modality, the explanation generation method, the application domain in which the explanation was evaluated, and the type of recommender system. Explanations might affect a user depending on their modalities or complexity, as different users might have other preferences. We could not observe any general patterns that certain explanation types are more likely to impact the results from our corpus. Possible reasons could be the small number of overall results or the necessity for a more fine-grained categorization of the explanation types. The impact of a user’s agreeableness on the perceived explanation quality, for example, was investigated by the same research team in two separate publications (Najafian et al., 2021a, b) focusing on privacy concerns regarding the information disclosed in the explanation. Both use text explanations in a group recommendation setting applied to the point-of-interest domain. The main difference in the explanations that arise from the articles is that the explanation in Najafian et al. (2021a) is static and discloses potentially private information, while the explanation in Najafian et al. (2021b) is presented in a chat interface that allows the user to interactively decide how much information is revealed to the group. Given the small data sample that we have for these results, applying a more fine-grained classification of explanation types would further increase the sparsity of the data.
@@
The articles of our corpus often did not disclose information on the underlying recommender system and explanation generation approach. One possible reason for this could be the focus of the articles. Describing the setup of a user study and the characteristics of the participants in sufficient detail requires space, which is usually restricted, especially in conference publications. This might have led to less detailed descriptions of the explanation method and underlying recommender systems. We furthermore identified a set of articles in which the recommendations and corresponding explanations were not generated by an actual recommender system but instead simulated for the user study (see (Shulner-Tal et al., 2022; Liao et al., 2022; Najafian et al., 2021a; Wibowo et al., 2018; Hernandez-Bocanegra et al., 2020; Coba et al., 2019), for example). These publications usually focus on investigating how different explanations impact the users and are often published in human-computer-interaction-centered venues such as CHI, UMAP, or IUI. It would be interesting to see in future work if these proposed explanations and results can be replicated with existing explanation methods. In the articles that provided information on the recommendation approach, we found a dominance of content-based and hybrid methods. This could indicate a trend toward providing information in the explanations that go beyond a user’s previous interactions or user-/item-similarity. Balog and Radlinski (2020), for example, showed that their baseline consisting of a short movie description is often perceived en par, if not better, than human-generated explanations for movie recommendations, indicating that providing additional information on the recommended items can be useful to achieve certain explanation goals.
@@
After categorizing the explanations, we still find different results for explanations with the same properties according to this classification, sometimes even from the same article. Kouki et al. (2020), for example, compared 11 different explanations and found that the visualization literacy of the participants impacts the quality perception of some explanations while others do not. Both explanations are visual explanations generated by an intrinsically explainable hybrid recommender system. While an effect was found for their Venn diagram explanation, it was not found for their cluster dendrogram. We leave investigating the detailed differences of such occurrences to determine the cause for these differences for future work.
@@ 5.5.1. Participant Recruiting
In terms of the recruitment process, we recommend a greater awareness of who is being recruited and whether this sample is representative of the target audience of the application domain. Furthermore, the shortcomings of the participant pool in crowdsourcing platforms need to be considered and addressed when opting for this type of recruiting. We suggest including a statement in the publication on what efforts were undertaken to ensure a representative participant sample with respect to the application domain. While the majority of application domains in our corpus have a low risk when a user follows the recommendations without understanding them, they usually target a diverse audience, which should be represented in the evaluation process. Domains such as the medical domain involve a higher risk compared to, e.g., POI recommendations. We, therefore, suggest paying special attention to the representativeness of the participant sample in these domains to make sure that the explanations are beneficial to all users. We particularly recommend paying attention to the participants’ location to not exclude users from the Global South.
@@ 5.5.5. Pre-Registration of User Studies
Pre-registering a study protocol is common practice in medicine and is increasingly promoted in social sciences to reduce reporting and publication bias (Munafò et al., 2017). We recommend fostering a discussion among the research community on how to establish an adoption of a pre-registration practice for human subject evaluation.

### v2

@@ 5.1. Who was recruited to evaluate recommender systems explanations?
The results regarding the education level of the participants show that the majority of the participants evaluating recommendation explanations are highly educated. Participants with at least a bachelor’s degree were the dominant group in our sample. The average participant has a higher education level than the average population.
@@ 5.3. Which user characteristics impact the effect of explanations in recommender systems?
In this survey, we looked at four potential factors, namely, the explanation modality, the explanation generation method, the application domain in which the explanation was evaluated, and the type of recommender system. Explanations might affect a user depending on their modalities or complexity, as different users might have other preferences. We could not observe any general patterns that certain explanation types are more likely to impact the results from our corpus. Possible reasons could be the small number of overall results or the necessity for a more fine-grained categorization of the explanation types. The impact of a user’s agreeableness on the perceived explanation quality, for example, was investigated by the same research team in two separate publications (Najafian et al., 2021a, b) focusing on privacy concerns regarding the information disclosed in the explanation. Both use text explanations in a group recommendation setting applied to the point-of-interest domain. The main difference in the explanations that arise from the articles is that the explanation in Najafian et al. (2021a) is static and discloses potentially private information, while the explanation in Najafian et al. (2021b) is presented in a chat interface that allows the user to interactively decide how much information is revealed to the group. Given the small data sample that we have for these results, applying a more fine-grained classification of explanation types would further increase the sparsity of the data.
@@
The articles of our corpus often did not disclose information on the underlying recommender system and explanation generation approach. One possible reason for this could be the focus of the articles. Describing the setup of a user study and the characteristics of the participants in sufficient detail requires space, which is usually restricted, especially in conference publications. This might have led to less detailed descriptions of the explanation method and underlying recommender systems. We furthermore identified a set of articles in which the recommendations and corresponding explanations were not generated by an actual recommender system but instead simulated for the user study (see (Shulner-Tal et al., 2022; Liao et al., 2022; Najafian et al., 2021a; Wibowo et al., 2018; Hernandez-Bocanegra et al., 2020; Coba et al., 2019), for example). These publications usually focus on investigating how different explanations impact the users and are often published in human-computer-interaction-centered venues such as CHI, UMAP, or IUI. In future work, it would be interesting to see if these proposed explanations and results can be replicated with existing explanation methods. In the articles that provided information on the recommendation approach, we found a dominance of content-based and hybrid methods. This could indicate a trend toward providing information in the explanations that go beyond a user’s previous interactions or user-/item-similarity. Balog and Radlinski (2020), for example, showed that their baseline consisting of a short movie description is often perceived en par, if not better, than human-generated explanations for movie recommendations, indicating that providing additional information on the recommended items can be useful to achieve certain explanation goals.
@@
After categorizing the explanations, we still find different results for explanations with the same properties according to this classification, sometimes even from the same article. Kouki et al. (2020), for example, compared 11 different explanations and found that the visualization literacy of the participants impacts the quality perception of some explanations while others do not. Both explanations are visual explanations generated by an intrinsically explainable hybrid recommender system. While an effect was found for their Venn diagram explanation, it was not found for their cluster dendrogram. We leave investigating the detailed differences of such occurrences to determine the cause of these differences for future work.
@@ 5.5.1. Participant Recruiting
In terms of the recruitment process, we recommend a greater awareness of who is being recruited and whether this sample is representative of the target audience of the application domain. Furthermore, the shortcomings of the participant pool in crowdsourcing platforms need to be considered and addressed when opting for this type of recruiting. We suggest including a statement in the publication on what efforts were undertaken to ensure a representative participant sample with respect to the application domain. While the majority of application domains in our corpus have a low risk when a user follows the recommendations without understanding them, they usually target a diverse audience, which should be represented in the evaluation process. Domains such as the medical domain involve a higher risk compared to, e.g., POI recommendations. We, therefore, suggest paying special attention to the representativeness of the participant sample in these domains to make sure that the explanations are beneficial to all users. We particularly recommend paying attention to the participants’ location so as to not exclude users from the Global South.
@@ 5.5.5. Pre-Registration of User Studies
Pre-registering a study protocol is common practice in medicine and is increasingly promoted in social sciences to reduce reporting and publication bias (Munafò et al., 2017). We recommend fostering a discussion among the research community on how to establish and adopt the pre-registration practice for human subject evaluation.
@@ 7. Conclusions and Future Work
Acknowledgements.
This work was partly supported by the Swiss National Science Foundation through project “MediaGraph” (contract no. 202125).